OpenStudy (anonymous):

Solve the equation by completing the square. 2x^2 + 9x + 1 = 0 x = (smaller value) x = (larger value)

OpenStudy (campbell_st):

ok... start by dividing each term by 2... what do you get..?

OpenStudy (anonymous):

x+9/2+1/2=0

OpenStudy (anonymous):

its is undefine?

OpenStudy (campbell_st):

not quite its \[x^2 + 4.5 x + 0.5 = 0\] next subtract 0.5 from both sides...what do you get..?

OpenStudy (anonymous):

x^2 + 4.5 = 0.5

OpenStudy (campbell_st):

a perfect square when expanded is \[(a + b)^2 = a^2 + 2ab + b^2\]

OpenStudy (campbell_st):

not quite its \[x^2 + 4.5x = -0.5\] does that make sense?

OpenStudy (anonymous):

iy eoulf look like x^2+4.5x+b2 = -0.5?

OpenStudy (anonymous):

it would*

OpenStudy (campbell_st):

if you are working in fraction the equation is \[x^2 + \frac{9x}{2} = - \frac{1}{2}\]

OpenStudy (campbell_st):

yes... so you need to find 4.5/2 =

OpenStudy (anonymous):

2.25

OpenStudy (campbell_st):

ok... to complete the square you now need to add (2.25)^2 to both sides...

OpenStudy (anonymous):

x^2 + 4.5 + 225 = - 0.5 +225

OpenStudy (anonymous):

2.25 i mean

OpenStudy (campbell_st):

so that you will have \[(x + 2.25)^2 = -0.5 + (2.25)^2\]

OpenStudy (campbell_st):

so the next step is \[(x + 2.25)^2 = 4.5625\] you just need to solve for x...

OpenStudy (anonymous):

square root?

OpenStudy (anonymous):

to take out the power of 2